OpenStudy (anonymous):

Three forces of 30 gms, 50 gms and 60 gms respectively act an angle of 120 degrees from each other. Find the resultant by rectangular-resolution (a)by making the 30-gm force lie on the axis. (b) by making the 60-gm force lie on the x-axis.

OpenStudy (anonymous):

have you computed it too? @Vincent-Lyon.Fr ? i'm not sure, but is the resultant 55.68?

OpenStudy (vincent-lyon.fr):

Before helping out with the calculation, let me make the following remark. Although it is not your fault, note that 'gms' in neither a valid symbol for mass nor force. symbol for gramme is 'g' : - no plural - no full stop; this is not an abbreviation but the the symbol of a mass-unit with a full mathematical value. It is not correct either to write: 'the 60-g mass' there should be no hyphen between 60 and g as '60 g' is in itself a quantity, in which 60 and g are factors in a multiplication. Reference is: point 5.3.3 of English SI-Brochure of BIPM http://www.bipm.org/utils/common/pdf/si_brochure_8_en.pdf

OpenStudy (anonymous):

ok..so the forces be 30 g, 50 g, and 60g...

OpenStudy (vincent-lyon.fr):

Resultant should be: F = \(\sqrt {700}\) gf = 26.46 gf * Fx = 30.cos0 + 50.cos120° + 60.cos240° Fy = 30.sin0 + 50.sin120° + 60.sin240° F = \(\sqrt{F_x^2+F_y^2}\) *gram-force (symbol gf) is outside the SI, but is nevertheless a valid unit of force.

OpenStudy (anonymous):

do you use rectangular resolution?

OpenStudy (vincent-lyon.fr):

I don't know what you mean by that, but I resolve each vector in the plane according to its orientation.